# Safe-SAIL: Towards a Fine-grained Safety Landscape of Large Language Models via Sparse Autoencoder Interpretation Framework

### v2

@@ Abstract
Increasing deployment of large language models (LLMs) in real-world applications raises significant safety concerns. Most existing safety research focuses on evaluating LLM outputs or specific safety tasks, limiting their ability to address broader, undefined risks. Sparse Autoencoders (SAEs) facilitate interpretability research to clarify model behavior by explaining single-meaning atomic features decomposed from entangled signals. However, prior applications on SAEs do not interpret features with fine-grained safety-related concepts, thus inadequately addressing safety-critical behaviors, such as generating toxic responses and violating safety regulations. For rigorous safety analysis, we must extract a rich and diverse set of safety-relevant features that effectively capture these high-risk behaviors, yet face two challenges: identifying SAEs with the greatest potential for generating safety concept-specific neurons, and the prohibitively high cost of detailed feature explanation. In this paper, we propose Safe-SAIL, a framework for interpreting SAE features within LLMs to advance mechanistic understanding in safety domains. Our approach systematically identifies SAE with best concept-specific interpretability, explains safety-related neurons, and introduces efficient strategies to scale up the interpretation process. We will release a comprehensive toolkit including SAE checkpoints and human-readable neuron explanations, which supports empirical analysis of safety risks to promote research on LLM safety.
1. Introduction
Increasing deployment of Large language models (LLMs) in critical applications raises significant safety concerns, including potential biases (Gallegos et al. 2024) and privacy breaches (Li et al. 2024). Previous studies have made great advances in safety-related areas from various perspectives. For instance, using classification models to measure the harmfulness of LLMs output (Hanu and Unitary team 2020; Lees et al. 2022), adversarial attack to identify vulnerabilities in LLMs (Schwinn et al. 2023), Chain-of-thought (CoT) monitoring to detect malicious behaviors in reasoning models (Baker et al. 2025), and latent space analysis for toxicity detection (Chacko et al. 2024; Xu et al. 2025).
Figure 1: Overview of safety-related SAE neuron database.
@@ 1. Introduction
Figure 2: Overview of the Safe-SAIL, which consists of three phases: SAE Training, Automated Interpretation, and Diagnose Toolkit. This framework trains sparse autoencoders with varying sparsity levels to select the most interpretable configuration, utilizes a large language model to explain neuron activations, and simulates query segments to calculate explanation confidence scores. Finally, the toolkit—including SAE checkpoints and a safety-tagged neuron database—is demonstrated through various case studies highlighting its applications in safety domains.
However, these approaches often focus on observable behaviors or pre-defined tasks. For example, one might first define a safety-related concept, such as hate, and then detect hateful content. Consequently, prior work is limited to specific tasks and cannot comprehensively address broader safety challenges. In contrast, we take an interpretability perspective by decomposing internal representations of LLMs, enabling us to obtain comprehensive safety concepts and identify undefined safety issues, such as cross-lingual alignment defects (see Section 4 for further discussion). We use Sparse Autoencoders (SAE) (Bricken et al. 2023; Cunningham et al. 2023) as our foundational tool: they factorize the entangled LLMs internal signals into a set of atomic features, without relying on supervision or pre-defined concepts. By interpreting the SAE features, we aim to uncover the underlying mechanisms that drive risk behaviors by providing a comprehensive fine-grained safety-related SAE neuron (Figure 1), which can be further used to diagnose, monitor, and potentially control undesired behaviors.
@@
Nevertheless, a significant gap remains between training SAE and providing human-aligned safety-related features, primarily arising from two aspects. Firstly, since it is computationally infeasible to generate and compare free-text explanations for every SAE configuration, selecting SAEs with the optimal configuration becomes challenging; thus, we need SAE evaluations prior to explanation generation. Most prior works employing SAEs (Lieberum et al. 2024; He et al. 2024) primarily evaluate SAEs using heuristic metrics, such as probing accuracy. They often lack evaluation of concept-specific interpretability of SAEs, that is, whether individual SAE features can differentiate nuanced concepts. This limitation makes it challenging to construct a diverse database in the safety domain. Second, generating human-readable explanations for SAE features and conducting evaluations (Bills et al. 2023; Choi et al. 2024; Paulo et al. 2024) require substantial resources. Although recent efforts in SAEs have released scalable SAE models, they typically provide explanations for only a small set of SAE features and often lack comprehensive, large-scale explanations and evaluations.
To address this gap, we propose Safe-SAIL, a Sparse Autoencoder Interpretation Framework for LLMs in safety domains. Our framework covers the entire process from SAE training, explanation generation and evaluation. Specifically, Safe-SAIL systematically selects the most effective SAEs to achieve the optimal diversity and interpretability. We build a bridge between SAE configurations and the diversity of neuron database in safety domains by new methods to evaluate concept-specific interpretability of SAEs. This provides practical guidance for selecting the SAE to produce neurons with optimal quantity and quality. Additionally, our approach replaces the traditional token-level simulation method with a more efficient segment-level strategy. We split the query into multiple segments and ask LRM to predict whether each segment is activated or not. This strategy reduces the simulation costs by 55% while maintaining satisfactory performance, making massive interpretation affordable. Hence, we finally generate human-understandable explanations for individual SAE neurons and provide comprehensive evaluations.
Moreover, we release an SAE toolkit based on the intermediate layers of Qwen 2.5-3B-Instruct (Qwen et al. 2025), which includes SAE checkpoints, explanations for individual safe SAE neurons and evaluations covering 2,059 SAE neurons across four major safety subdomains: pornography, politics, violence, and terror (Figure 1). Our toolkit enables fine-grained analysis of internal safety mechanisms and facilitates monitoring of LLMs’ risk behaviors, thereby supporting broader adoption and future research. Based on the toolkit, we conduct some empirical analyses on pornographic concepts, demonstrating the potential of Safe-SAIL for risk identification in LLMs. Our investigation yields several intriguing findings, including insights into how LLMs encode specific real-world risk entities and handle safety-critical concepts related to sexually explicit content.
The main contributions of this work include:
  • We propose Safe-SAIL, a framework for interpreting SAEs in safety domains. It offers a perspective by decomposing LLMs’ internal representations to identify comprehensive and undefined safety features. To support further research and practical applications, we release an open-source toolkit based on Safe-SAIL, including 2,059 neurons across four major safety subdomains, and analytical utilities.
  • We improve the efficiency of SAE interpreting by introducing two key strategies: a concept-specific interpretability evaluation method that enables optimal SAE model selection before explanation generation, and a segment-level simulation approach that significantly reduces computational overhead, making massive evaluation affordable.
@@
  • Based on our toolkit, we conduct empirical analyses on pornographic concepts, demonstrating the potential of Safe-SAIL for risk identification in LLMs. And we offer insights on how LLMs encode specific real-world risk entities and handle safety-critical concepts across layers.
2. Framework
In this section, we introduce the three components of the Safe-SAIL, as illustrated in Figure 2. The first component is SAE training and evaluation, which focuses on training and selecting an SAE that produces the most interpretable features in safety domains. The second component is automated interpretation, including free-text explanation generation, and evaluations of explanations, thus facilitating human understanding. Applying this framework, we obtain a new set of toolkit to carry out safety analysis.
2.1 Sparse Autoencoders
@@ 2.1.1 Training
We utilize Sparse Autoencoders (SAEs) that incorporate the TopKReLU activation function(Gao et al. 2024) to control sparsity levels in the encoded representation. Given an input signal $x\in\mathbb{R}^{D}$, typically derived from the output of Multilayer Perceptrons (MLPs) or Residual Streams, the TopKReLU activation function selects the top-k activated features during the encoding transformation. Details of SAE training and TopKReLU are in Appendix.
@@ 2.1.2 Enhanced Evaluation Metrics
To interpret neurons related to concepts in safety domains, our primary concern is whether the SAE has greater potential to differentiate and extract more nuanced atomic concepts within that domain. However, the time and computational costs associated with training the SAE and interpreting all its neurons can be exceedingly high. Therefore, we seek a metric that can predict the actual number of neurons related to safety concepts that will be effectively explained after completing the SAE interpretation. We construct evaluation data using a method Concept Contrastive Query Pairs to illustrate the boundaries of the presence or absence of concepts. Additionally, we design two metrics, $L_{0,t}$ and $I_{CDF}$, to assess the differentiation of concepts among different SAEs.
Concept Contrastive Query Pairs We prepare a dataset consisting of queries categorized under various safety domains. For each query related to a specific concept theme, we design prompts that instruct LLMs to generate a paired query that omits this particular concept while preserving the other linguistic elements as closely as possible.
Metrics For each concept domain with $n$ pairs, we collect the delta frequency freqfreq of each neuron that activates on concept query while not on the de-concept paired one. $Q_{C}$ and $Q_{D}$ denote whether this neuron activates on concept query or corresponding de-concept one.
$freq=\frac{\sum_{i=0}^{n-1}Q_{C,i}(1-Q_{D,i})}{n},\ Q_{C,i},Q_{D,i}\in\{0,1\}$ (1)
For each concept theme, all neurons on SAE could be represented by first a distribution frequency function and second a cumulative distribution frequency function denoted as:
@@
We describe the interpretability of an SAE from the following aspects.
  • $L_{0,t}$ discovers absolute number of distinguishable neurons in a specific safety domain. The value varies by threshold, which can differ across domains; for this analysis, a threshold of 0.25 is employed, as it has been empirically observed to enable distinguishability between all SAEs compared.
    $L_{0,t}=\sum(freq>t)\\$ (4)
  • $I_{CDF}$ represents the expected delta frequencies of all neurons in the set, reflecting the overall distinguishability of the entire SAE in relation to a specific thematic concept; it allows for intuitive comparison by visualizing the area under the curve in the CDF plot.
@@
2.2 Efficient Neuron Interpretation
2.2.1 Safety-Neuron Filtering
The overall cost of interpretation is high, and the number of safety-related neurons is relatively small compared to the total. Therefore, we first employ a filtering method to obtain candidate neurons. Using the method of Concept Contrastive Query Pairs, we construct concept de-concept pairs based on a more fine-grained classification of safety data. We observe the activation patterns of neurons across each subclass; a neuron that is associated with a specific safety concept should exhibit a noticeable difference in activation distribution between the concept and de-concept sets. However, considering that the concept scope corresponding to a given neuron may be narrower than our classification definitions, we should establish a low recall threshold and a high precision threshold when calculating accuracy and recall using the following expressions:
$Precision=\frac{\sum Q_{C}}{\sum Q_{C}+\sum Q_{D}},Recall=\frac{\sum Q_{C}}{n}$ (6)
@@ 2.2.1 Safety-Neuron Filtering
Precision refers to the ratio of activated concept queries to the total activated queries. Recall indicates the ratio of activated concept queries to the total concept queries.
2.2.2 Explanation
@@
We adopt the standard practice for generating neuron explanations: neuron activations are generated through SAE inference on an explanation dataset rich in safety-related content. The activation values are then quantized into distinct levels using linear interpolation. For each level, samples are selected to construct a prompt that instructs a large reasoning model (LRM) to generate a text explanation for the corresponding neuron.
2.2.3 Segment-level Simulation & Scoring
One of the most common methods for evaluating explanations is simulation. In traditional simulation, an LLM is used to predict the activations of each token in a query, given both the neuron explanation and the tokenized query. The simulation score, referred to as the CorrScore, is then calculated as the Pearson correlation coefficient between the simulated activations and actual token activations after inference.
However, high-quality simulations always require high computational resources. To optimize the simulation process, we first use LRM to infer activation values at multiple positions in a single call, rather than predicting the activation for each token in separate forward passes (Bills et al. 2023). However, obtaining a reliable simulation score still requires sampling a substantial number of query, which results in significant computational overhead. To address this, we split each query into several segments and instruct LRM to predict whether each segment will be activated by the neuron. Larger internal segmentation leads to lower computation but poor scoring performance.
2.3 Toolkit
With the safety neuron database constructed, we provide an interactive tool and a feature map that allow researchers to explore which safety-relevant neurons are activated by arbitrary inputs and query their semantic interpretations. The database also serves as the foundation for two key insights related to knowledge detection and interpretable inference trajectories. A detailed discussion can be found in Section 4.
Table 1: Compare SAEs trained with different settings from reconstruction and interpretability. We also explained neurons in these SAEs to construct a safety-related neuron database to illustrate how SAE configuration influences neuron explanation quantity and quality. Details of metrics are included in appendix.
@@ 2.3 Toolkit
Location TopK $R_{alive}\uparrow$ Reconstruction Interpretability Neuron Database
@@
  • $R_{alive}$: Percentage of neurons triggered during inference. $L_{2}$: MSE between SAE input and reconstructed output. $\delta L_{NTP}$: Difference in next token prediction loss. $L_{0,t=0.25}$: Number of neurons whose freqfreq larger than 0.25.
  • $I_{CDF}$: Expected value of freqfreq across all neurons in SAE. NN: Number of concept-specific neurons. CorrScoreCorrScore: Average correlation score of all safety-related neurons. SpScoreSpScore: Average superposition score of all safety-related neurons.
Figure 3: Cumulative distribution frequency curve of SAEs trained with different settings.
Figure 4: Neurons related to concept of adult content from neuron databases derived from different SAE checkpoints. The distribution illustration is based on distance between text embeddings of neuron explanations.
Figure 5: Comparison of various interpretability metrics against ground truth across different sparsity levels $L_{0}$ and multiple safety domains. (a) Ground truth showing the number of concept-specific neurons. (b) Our proposed metrics: $I_{CDF}$ and $L_{0,t=0.25}$, demonstrating trends closely aligned with the ground truth. (c) 1d-Probe cross entropy loss varies in different safety-domains. (d) $k$-Sparse Probing performance (with kk=1,3,5,20) depends largely on $k$.
3. Experiments
In this section, we investigate the impact of selecting optimal parameters within each stage of our proposed framework. Our primary goal is to demonstrate how these parameter choices lead to improved results and reduced costs.

### v3

@@ Abstract
Sparse autoencoders (SAEs) enable interpretability research by decomposing entangled model activations into monosemantic features. However, under what circumstances SAEs derive most fine-grained latent features for safety—a low-frequency concept domain—remains unexplored. Two key challenges exist: identifying SAEs with the greatest potential for generating safety domain-specific features, and the prohibitively high cost of detailed feature explanation. In this paper, we propose Safe-SAIL, a unified framework for interpreting SAE features in safety-critical domains to advance mechanistic understanding of large language models. Safe-SAIL introduces a pre-explanation evaluation metric to efficiently identify SAEs with strong safety domain-specific interpretability, and reduces interpretation cost by 55% through a segment-level simulation strategy. Building on Safe-SAIL, we train a comprehensive suite of SAEs with human-readable explanations and systematic evaluations for 1,758 safety-related features spanning four domains: pornography, politics, violence, and terror. Using this resource, we conduct empirical analyses and provide insights on the effectiveness of Safe-SAIL for risk feature identification and how safety-critical entities and concepts are encoded across model layers. All models, explanations, and tools are publicly released in our open-source toolkit [^1] https://github.com/Alibaba-AAIG/Safe-SAIL and companion product [^2] https://modelscope.cn/studios/Alibaba-AAIG/Safe-SAIL/summary.
Safe-SAIL: Towards a Fine-grained Safety Landscape of Large Language Models via Sparse Autoencoder Interpretation Framework
Warning: this paper contains data, prompts, and model outputs that are offensive in nature.
Jiaqi Weng [^1] footnotemark: 1,  Han Zheng211footnotemark: 1,  Hanyu Zhang1,  Ej Zhou3,  Qinqin He1,  Jialing Tao122footnotemark: 2, Hui Xue1,  Zhixuan Chu [^2] footnotemark: 2,  Xiting Wang4 1Alibaba Group  2The State Key Laboratory of Blockchain and Data Security, Zhejiang University   3Language Technology Lab, University of Cambridge  4Renmin University of China {wengjiaqi.wjq, jialing.tjl}@alibaba-inc.com, {h.zheng}@zju.edu.cn
1 Introduction
Increasing deployment of large language models (LLMs) in critical applications raises significant safety concerns Gallegos et al. (2024); Li et al. (2024). Previous studies have made advances in LLM safety from various perspectives Hanu and Unitary team (2020); Lees et al. (2022); Schwinn et al. (2023); Baker et al. (2025); Chacko et al. (2024); Xu et al. (2025); however, these approaches often focus on observable behaviors or pre-defined tasks, leaving them task-bound and blind to wider, unseen risks. A recent line of work approaches safety through interpretability methods, aiming to decompose and analyze internal representations of LLMs to reveal latent mechanisms and safety-relevant features Chen et al. (2024); Xu et al. (2024); Arditi et al. (2024); Zhao et al. (2025); In particular, sparse autoencoders (SAEs) Bricken et al. (2023); Cunningham et al. (2023) have drew on much attention: they factorize the entangled internal signals into a set of atomic features, without relying on supervision or pre-defined concepts. Its structured feature space would enable us to uncover the underlying mechanisms that drive risk behaviors (Figure 1), which can be further used to diagnose, monitor, and potentially control undesired behaviors.
Figure 1: Overview of our trained safety‑related SAE feature-base. The feature-base covers four safety domains; here we show the pornography domain as an example. In total, 309 features from SAEs trained on quarter layers (0, 8, 17, 26, 35) are interpreted as related to pornography. The chart on the right lists the top five sub‑categories of these features.
@@ 1 Introduction
Figure 2: Overview of our framework Safe-SAIL, which consists of three phases: SAE Training, Automated Interpretation, and Diagnose Toolkit. This framework trains sparse autoencoders with varying sparsity levels to select the most interpretable configuration; utilizes LRM to explain feature activations and simulates query segments to calculate explanation confidence scores; and facilitate various case studies with the acquired SAE checkpoints and a tagged feature database in safety domains.
@@
Nevertheless, a significant gap remains between training SAEs and delivering human-aligned safety-related features, primarily due to two challenges. First, generating and comparing free-text explanations for every SAE configuration is computationally infeasible, making it difficult to identify optimal configurations. Efficient pre-explanation evaluation metrics are therefore essential. Most prior works employing SAEs Lieberum et al. (2024); He et al. (2024) primarily evaluate SAEs using heuristic metrics, such as probing accuracy. They often lack evaluation of domain-specific interpretability of SAEs, that is, whether individual SAE features can differentiate nuanced concepts. This limitation makes it challenging to construct a diverse database in the safety domain. Second, generating human-readable explanations for SAE features and conducting evaluations Bills et al. (2023); Choi et al. (2024); Paulo et al. (2024) require substantial resources. Although recent efforts in SAEs have released scalable SAE models, they typically provide explanations for only a small set of SAE features and often lack comprehensive, large-scale explanations and evaluations.
To address this gap, we propose Safe-SAIL, a Sparse Autoencoder Interpretation Framework for LLMs in safety domains. Our framework covers the entire process from SAE training, explanation generation and evaluation, as illustrated in Figure 2.
First, Safe-SAIL systematically selects SAEs with optimal domain-specific interpretability. We introduce new evaluation metrics that quantify how well SAE configurations differentiate safety concepts (§ 2.1.2). This provides practical guidance for selecting SAEs that produce features with optimal quantity and quality for safety analysis. Second, to reduce the prohibitive cost of explanation evaluation, we replace traditional token-level simulation with a segment-level strategy. Specifically, we split each query into $n$ segments and use a large reasoning model (LRM) to predict binary activation status (activated/not activated) for each segment. As shown in § 2.2, this approach reduces simulation costs by 55% while maintaining high correlation with token-level results, making large-scale interpretation feasible. Third, we release an open-source SAE toolkit that provides an interactive interface and a feature map for exploring safety-relevant features activated by arbitrary inputs and querying their semantic interpretations (§ 2.3). Building upon the Safe-SAIL framework, we train a comprehensive suite of SAEs on Qwen2.5-3B-Instruct Qwen et al. (2025) and generate human-readable explanations for safety-related features across four subdomains: pornography, politics, violence, and terror. Using this resource, we conduct empirical analyses on pornographic concepts, yielding insights into: (1) how LLMs encode specific real-world risk entities (§ 4.1), and (2) how they handle safety-critical concepts related to sexually explicit content (§ 4.2).
The contributions of this work are:
  • We introduce Safe-SAIL, a novel framework for interpreting SAEs in safety-critical domains. Safe-SAIL enables efficient identification of safety domain-specific features through a new pre-explanation evaluation metric and reduces interpretation cost by 55% via segment-level simulation. The code and full implementation are publicly released.
  • Building on Safe-SAIL, we release a suite of SAEs trained on Qwen2.5-3B-Instruct, accompanied by human-readable explanations and evaluations for 1,758 safety-related features across four subdomains: pornography, politics, violence, and terror.
@@
  • Based on this resource, we conduct empirical analyses on pornographic concepts, demonstrating the potential of Safe-SAIL for risk identification in LLMs. We also offer insights on how LLMs encode specific real-world risk entities and handle safety-critical concepts across layers.
2 Framework
This section details the three principal components of the Safe-SAIL framework, depicted in Figure 2.
2.1 SAE Training
@@ 2.1.1 Training
We employ SAEs to decompose dense internal activations of LLMs into a higher-dimensional, sparse feature representation. Given an input signal $x\in\mathbb{R}^{D}$, typically derived from the output of multilayer perceptrons (MLPs) or residual streams, an SAE encodes $x$ into a sparse latent code $z$, where only a small subset of features are active, and then reconstructs an approximation $\hat{x}$ from $z$. For our training setup, details can be found in § 3.1.1.
@@ 2.1.2 Enhanced Evaluation Metrics
Training and interpreting all SAE features is prohibitively expensive, therefore, we seek a metric that can predict the number of valid features after completing the SAE interpretation. We construct evaluation data using Concept Contrastive Query Pairs, which consist of paired examples that contrast the presence and absence of a target concept. We design two metrics, $L_{0,t}$ and $I_{CDF}$, to assess the differentiation of concepts among different SAEs.
Concept Contrastive Query Pairs We prepare a dataset consisting of queries categorized under various safety domains. For each query related to a specific concept theme, we design prompts that instruct LLMs to generate a paired query that omits this particular concept while preserving the other linguistic elements as closely as possible Bohacek et al. (2025).
Metrics For each concept domain with $n$ pairs, we collect the delta frequency freqfreq of each latent ($freq_{k}$ for latent $k$) that activates on concept query while not on the de-concept paired one. $Q_{C}$ and $Q_{D}$ denote whether this feature activates on concept query or corresponding de-concept one.
$freq_{k}=\frac{\sum_{i=0}^{n-1}Q_{C,i}(1-Q_{D,i})}{n},$ (1)
where $Q_{C,i},Q_{D,i}\in\{0,1\}$.
For each concept theme, all latents on SAE could be represented by first a distribution frequency function and second a cumulative distribution frequency (CDF) function denoted as:
@@
We describe the interpretability of an SAE from the following aspects (see Appendix B for details).
  • $L_{0,t}$ discovers the absolute number of distinguishable latent features in a specific domain. The value varies by the chosen threshold $t$, which can be flexibly adjusted based on the research context and target domain.
    $L_{0,t}=\sum_{k=0}^{M-1}\begin{cases}1,&\text{if }freq_{k}>t\\ 0,&\text{if }freq_{k}\leq t\end{cases}\\$ (4)
  • $I_{CDF}$ represents the expected delta frequencies of all features in the set, reflecting the overall distinguishability of the entire SAE in relation to a specific thematic concept; it allows for intuitive comparison by visualizing the area under the curve in the CDF plot.
@@
2.2 Automated Interpretation
Safety-Feature Filtering
We employ a filtering method to identify safety-relevant candidate features. Specifically, we construct Concept Contrastive Query Pairs based on evaluation data and examine feature activation patterns across subclasses. A feature associated with a specific safety concept should exhibit significant differences in activation distribution between concept and de-concept sets.
$Precision=\frac{\sum Q_{C}}{\sum Q_{C}+\sum Q_{D}},\quad Recall=\frac{\sum Q_{C}}{n}$ (6)
@@ 2.2 Automated Interpretation
Precision refers to the ratio of activated concept queries to the total activated queries. Recall indicates the ratio of activated concept queries to the total concept queries. To filter concept sensitive features, we set two thresholds ($t_{p}$ and $t_{r}$) for precision and recall respectively. Features are selected if they satisfy $Precision>t_{p}$ and $Recall>t_{r}$ simultaneously. Details on feature filtering are in Appendix B.
Explanation
@@
We adopt the standard practice Bills et al. (2023); Paulo et al. (2024) for generating feature explanations: feature activations are generated through SAE inference on a customized explanation dataset. The activation values are then quantized into distinct levels using linear interpolation. For each level, samples are selected to construct a prompt that instructs a large reasoning model (LRM) to generate a text explanation for the corresponding feature.
Segment-level Simulation & Scoring
Previous works Bills et al. (2023); Paulo et al. (2024) evaluate explanations with simulation, where an LLM is used to predict the activations of each token in a query, given both the neuron explanation and the tokenized query. The simulation score, referred to as the CorrScore, is then calculated as the Pearson correlation coefficient between the simulated activations and actual token activations after inference. Apparently, high-quality simulations require high computational resources. To optimize the simulation process, we first use LRM to predict activation value at each token in a query in one call, rather than predicting the activation for each token in separate forward passes Bills et al. (2023). However, we find that predicting all tokens in a single call generates excessively long LRM responses, which causes substantial computational overhead. To address this, we split each query into $n$ segments and instruct LRM to predict neuron activations on each segment.
2.3 Diagnose Toolkit
With the feature database constructed, we provide an interactive diagnostic toolkit comprising an interactive tool and a feature map. The interactive tool allows researchers to input an arbitrary query and visualize, for each token position, those safety-relevant features with the highest activation values. For each activated feature, the tool displays its human-readable semantic explanation and the associated correlation score. The feature map offers a global view of the safety feature landscape. It projects all annotated features into a 2D space where the spatial distance between any two points reflects their semantic similarity. This map enables intuitive navigation, clustering analysis, and the discovery of relationships among safety concepts within the model’s latent space.
@@ 2.3 Diagnose Toolkit
Location TopK $R_{alive}\uparrow$ Reconstruction Interpretability Feature Database
@@
  • $R_{alive}$: Percentage of features triggered during inference.
  • $L_{2}$: MSE between SAE input and reconstructed output.
  • $\delta L_{NTP}$: Difference in next token prediction loss.
  • $L_{0,t=0.25}$: Number of features whose freqfreq larger than 0.25.
  • $I_{CDF}$: Expected value of freqfreq across all features in SAE.
  • NN: Number of safety domain features.
  • CorrScoreCorrScore: Average correlation score of all related features.
  • SpScoreSpScore: Average superposition score of all related features.
Table 1: Comparing SAEs trained with different settings from reconstruction and interpretability. We also explain features in these SAEs to construct a safety-related feature database to illustrate how SAE configuration influences feature explanation quantity and quality. Details of metrics are included in Appendix B.
Figure 3: CDF curve of SAEs trained with different settings.
Figure 4: Features on concept of adult content from different SAE checkpoints. The distribution illustration is based on distance between text embeddings of feature explanations.
Figure 5: Comparison of various interpretability metrics against ground truth across different sparsity levels $L_{0}$ and multiple safety domains. (a) Ground truth showing the number of domain-specific features. (b) Our proposed metrics: $I_{CDF}$ and $L_{0,t=0.25}$, demonstrating trends closely aligned with the ground truth. (c) 1d-Probe cross entropy loss varies in different safety-domains. (d) $k$-Sparse Probing performance (with kk=1,3,5,20) depends largely on $k$.
3 Experiments
In this section, we investigate the impact of selecting optimal parameters within each stage of our proposed framework on Qwen2.5-3B-Instruct, specifically within the context of safety domains. Our primary goal is to demonstrate how these parameter choices lead to improved results and reduced computational costs. We use the identified best parameter settings to generate a series of SAE checkpoints, which not only underpin our subsequent empirical analyses but are also publicly released as part of our contribution.